OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each. f(x) = 5(x + 8)^2(x - 8)^3

OpenStudy (solomonzelman):

When you have \(\large\color{blue}{ a \times b=0 }\), then either \(\large\color{black}{ a =0 }\) or, \(\large\color{black}{ b=0 }\).

OpenStudy (solomonzelman):

This is called a zero product property.

OpenStudy (solomonzelman):

Here, you got a similar thing, (because you want to find the zeros of the function, you set \(\large\color{black}{ f(x) =0 }\) ) \(\large\color{blue}{ 0= 5(x + 8)^2(x - 8)^3 }\) So, either \(\large\color{black}{ 0= (x + 8)^2 }\) or \(\large\color{black}{ 0= (x - 8)^3 }\)

OpenStudy (anonymous):

@SolomonZelman the answer choices are A) 4, multiplicity 1; -8, multiplicity 3; 8, multiplicity 3 B) -8, multiplicity 2; 8, multiplicity 3 C) -8, multiplicity 3; 8, multiplicity 2 D) 4, multiplicity 1; 8, multiplicity 1; -8, multiplicity 1 so would it be c?